LHC 2011 - latest news

Week 7: Machine coordinators: Jorg Wenninger - Jan Uythoven

*************************************************** During the weekend: morning meeting at 9:30 ************

Main aims of the week:

LHC coordination vistar

Commissioning outline



Link to report from morning meeting

Monday 21st February
  • Official start of beams.
08:30 meeting
Sunday 20th February
  • RF capture setup.
  • Beta-beat measurements, BI setup, BLM system checks.
  • Q and Q' decay overnight.
09:30 meeting
Saturday 19th February
  • Access in the morning for QPS and EPC.
  • Evening & night:
    • MKI synchronization with beam on TEDs.
    • B2 threading and splashes. Beam circulating at 22:30.
    • B1 threading and splashes. Beam circulating at 01:00.
    • Basic checks at 450 GeV.
09:30 meeting - checkout
Friday 18th February
  • Access in the morning, handshake test.
  • BETS tests at 450 GeV completed.
  • Evening: move through nominal sequence.
    • Pre-cycle: OK for all circuits except RCD (dI/dt issue).
    • Machine to injection (OK except RCD), arm BIS loop. Dump tests.
    • Ramp to 3.5 TeV with BIS loop armed. Checked PM disable not possible above 450 GeV.
    • Squeeze to 1.5/10/1.5/3 m (sep/Xing to 0) OK. Beta* information correct. Checked MDI flag in beam dump mode. Beam dump at 3.5 TeV.
    • Overall pretty good, lot's of tasks to tune/adapt.
  • 01:00 Heat run. Trip of S12.
  • Beam to TI2 and TI8, test of trajectory dependence on SPS SC composition.
16:00 meeting - checkout
Thursday 17th February
  • 13:00-22:00: ring closed, close BIS loop for MKI/LBDS.
  • 16:00: B2 permit loop closed, problems on B1 loop (BIS-LBDS link). Access, repair optical receiver on TSU.
  • 20:00 Both BIS loops closed. Problems with link from LBDS to MKI. Ramp to 3.5 TeV (BETS PCs) and programmed dump test.
  • 20:00-24:00: TI2 and TI8 transfer lines tests

8:30 meeting HWC, checkout


Wednesday 16th February
  • 16:00-24:00: ring closed, close BIS loop for MKI/LBDS.
  • Transfer line tests:
    • 15:00-02:00: beam to downstream TI8 TED
    • 17:00-02:00: beam to downstream TI2 TED

8:30 meeting HWC checkout

Tuesday 15h February
  • SPS:
    • Probe bunch ready in the ring.
    • Extraction setting up. Beam to TT40 TED.
  • 16:00:2400: Ring closed.
    • Energy tracking test to 3.5 TeV.

8:30 meeting

Monday 14th February

  • Morning: handshake test and SMP (with simulated input E + b*)
  • Afternoon: SMP connected to real energy inputs.
  • 18:00-24:00: ring closed, attempt to close BIS loop for MKI/LBDS.


Status of Cold Checkout